OpenStudy (anonymous):

A consumer survey was taken in a town with 18,000 homes. Of the 360 homes surveyed, 144 had computers. On the basis of this survey, estimate the number of homes in the town that have computers.

OpenStudy (anonymous):

what you mean by that

OpenStudy (ash2326):

let's assume that the average no. of homes having computer is the same as that found from the survey let's calculate the percentage of homes having computer (which took part in the survey) 144/360*100=40% now we assume this is also the average for all the homes so homes having computer in whole town=18000*40%=7200

OpenStudy (mr.math):

\(\frac{144}{360}\times 18,000=7200\text{ homes }\).

OpenStudy (anonymous):

u need help